OpenStudy (anonymous):

6 over a-8 subtracted by 6 over 8-a

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

Well for the 6/8-a, it can also be written as 6/-a+8.

OpenStudy (anonymous):

okayy then what? im confused:(

OpenStudy (anonymous):

Then you can factor so it looks really similar to the other one. 6/-a+8 = 6/-(a-8)... So the only difference is..?

OpenStudy (anonymous):

so 6/(8-a)=-(6/(a-8)) therefore the answer is (6/(a-8))-(6/(8-a))=(6/(a-8))+(6/(a-8))=12/(a-8)

OpenStudy (aravindg):

guys giv amedal for me i will giv in return